# Strategic Demand-Planning in Wireless Networks: Can Generative-AI Save Spectrum and Energy?

### v1

@@ III-D Towards Resilient Networking: Managing Disasters and Unusual Gatherings
According to the discussions in Sections II and III, it is evident that GenAI will be an indispensable part of the future of wireless networks. On the flip side, implementing GenAI in wireless networks for demand-planning, particularly demand-shaping, purposes has its own challenges, and in the following paragraphs, we highlight some of the key challenges.
IV-A Hardware Bottlenecks
Demand-shaping by GenAI can be implemented at user terminals or BSs, considering the objective of the implementation. Although BSs are equipped with a relatively more powerful hardware architecture that facilitates the implementation of AI algorithms, arming user equipment with GenAI might be a bottleneck in realizing GenAI-based demand-shaping at the device level. Albeit the attempts in industry to make AI available on user terminals, demand-shaping through GenAI requires high computation and processing capabilities to compress the data, boosting the energy consumption and heat generation that subsequently affects the current and long-term performance of the device. Hence, it is vital to consider the user terminal hardware bottlenecks for implementing GenAI demand-shaping in future wireless networks.
IV-B Cyber Security
Ensuring the security of data and GenAI models is crucial for the feasibility and applicability of the demand-shaping concept. GenAI models require a robust system capable of storing and processing large datasets to shape user demand. Managing the vast volume of data generated by GenAI presents a significant challenge. During the demand-shaping process, there is a risk of malicious interventions aimed at disrupting the system. Additionally, data may be accessed, monitored, or manipulated by malicious parties, posing serious privacy concerns. Such abuses not only threaten system functionality but also highlight the need for robust security measures. Consequently, a strong and expandable infrastructure is needed to securely store, access, and process this data. Protecting the data from unauthorized access and implementing appropriate security measures is challenging, especially when it includes confidential information. The research highlighted in [15] underscores the importance of addressing security vulnerabilities, emphasizing that AI can potentially increase cyber attacks. Therefore, employing robust encryption methods is vital to sustain privacy during data transmission and storage.
IV-C Information Loss
@@
A common flaw in both voice and text data shaping is the oversimplification of information. Although GenAI algorithms generally perform with good accuracy, there is still a risk of losing contextual information during the demand-shaping process. This can result in the omission of important details and nontrivial nuances, leading to incomplete information for the receiver. Therefore, it is crucial to have efficient mechanisms to prevent information loss. In scenarios where information loss is unavoidable, data rescheduling is a potential replacement solution for demand-shaping.
IV-D Ethical Perspective
Implementation of demand-shaping through GenAI models will impose several ethical concerns. According to the 2024 AI Index report [15], a research study among AI manufacturers and users, highlights the absence of standardized evaluations for AI models, the increasing prevalence of AI-related incidents, and concerns about transparency and security. Some of the prominent ethical concerns of implementing GenAI are mentioned in the following.
IV-D1 Consistency and Reliability
@@
GenAI models can sometimes produce inconsistent outputs for the same content, leading to unreliable or unpredictable results. This is especially critical in situations such as disasters, where models need to operate smoothly and quickly. Therefore, it is essential to develop and continuously improve validation methods to enhance the reliability of GenAI models.
IV-D2 Data Privacy, Transparency and Accountability
When processing user demands, the security and confidentiality of data are of high importance. GenAI will analyze and potentially store the user data to shape the demands. However, monitoring or storing users’ sensitive data may lead to privacy and security risks. Additionally, the black box nature of GenAI, with its lack of transparency about how algorithms work, can make it difficult for users to trust the process. Furthermore, the absence of accountability can hinder the correction of potential errors or misunderstandings in the models.
@@ V Conclusion
Berk Çiloğlu (cilogluberk@gmail.com) completed his B.Sc. degree in Electrical and Electronics Engineering at Ankara Yıldırım Beyazıt University, Ankara, Türkiye in 2023. His research focuses on applications of 6G and beyond wireless communication technologies, and non-terrestrial networks.
Görkem Berkay Koç (gorkembrkoc@gmail.com) received a B.Sc. degree in Electrical and Electronics Engineering at Ankara Yıldırım Beyazıt University, Ankara, Türkiye, in 2023. His research interests focus on optimizing wireless network performance, specifically non-terrestrial networks, and visible light communications.
Afsoon Alidadi Shamsabadi (afsoonalidadishamsa@sce.carleton.ca) is a Ph.D. Candidate in the Non-Terrestrial Networks (NTN) Lab at Carleton University, Canada. Her research focuses on the innovative integration of High Altitude Platform Stations (HAPS) with terrestrial wireless networks.
Metin Ozturk (metin.ozturk@aybu.edu.tr) is currently a Visiting Professor at Carleton University, Canada, and an Assistant Professor at Ankara Yıldırım Beyazıt University, Turkiye. His current research includes intelligent networking for wireless communications, with a particular focus on non-terrestrial networks (NTN) and sustainability.
Halim Yanikomeroglu (halim@sce.carleton.ca) is a Chancellor’s Professor in the Department of Systems and Computer Engineering at Carleton University, Ottawa, Canada, leading the Non-Terrestrial Networks (NTN) Lab. His research group focuses on wireless access architecture for the 2030s and 2040s, with a particular emphasis on NTN.

### v2

@@ III-D Towards Resilient Networking: Managing Disasters and Unusual Gatherings
III-E Use-Cases and Business Implications
The usage scenarios, enabled by the proposed GenAI-based demand shaping framework, have the potential to facilitate new use-cases with significant business implications. The aforementioned usage scenarios support network operators in managing the network with the existing available spectrum and energy resources. Considering that the spectrum is an expensive asset for the operators, this significantly supports them in saving the cost of acquiring additional spectrum. Additionally, the remnant spectrum that is saved with the proposed approaches can be repurposed, such as leasing it to other operators. In this way, primary operators can increase their profits through a new revenue stream from spectrum leasing, while secondary operators gain more opportunities to enhance the quality of service (QoS) for their users.
Furthermore, the energy savings from BS switching are reflected in reduced operational costs, making the businesses of operators more economically sustainable. The more BSs that can be switched off, the greater the potential impact. GenAI-based demand shaping plays a crucial role here, as it enables the switching off of more BSs by reducing their load. In addition, the usage scenario described in Section III-D is a critical use case where the proposed methodology can save lives during disaster situations. Moreover, it plays a vital role in unusual gatherings by improving or maintaining user satisfaction, a key factor in business sustainability.
@@ IV Challenges and Open Issues
According to the discussions in Sections II and III, it is evident that GenAI will be an indispensable part of the future of wireless networks. On the flip side, implementing GenAI in wireless networks has its own challenges, and in the following paragraphs, we highlight some of the key challenges.
IV-A Hardware Bottlenecks and Data Content Access
Although BSs are equipped with a relatively more powerful hardware architecture, arming user equipment with GenAI might be a bottleneck in realizing GenAI-based demand-shaping at the device level. Furthermore, the BS faces challenges in applying GenAI due to its lack of access to application-layer data. Edge computing offers a solution by handling content analysis and AI tasks at nearby nodes, reducing the processing burden on the BS.
IV-B Regulatory Policies and Carbon Emission
Regulatory policies, such as net neutrality, aim to maintain a fair data service for all the data, regardless of its source, content, or destination. Demand-planning through GenAI can theoretically be designed to support net neutrality by ensuring that all users with the same service-level agreements (SLAs) receive equal treatment. At the same time, the GenAI should not prioritize traffic unfairly or violate these agreements. In any scenario, prioritizing demand planning during emergencies can be justified ethically and legally to protect public safety, provided such actions are temporary. Demand-shaping through GenAI requires high computation and processing capabilities to compress the data, boosting the energy consumption, and therefore the carbon emission which is a critical issue. GenAI algorithms are applied in demand shaping for good reasons, but the increased carbon emissions resulting from these computationally demanding algorithms should not overshadow the benefits achieved in demand shaping.
IV-C Cyber Security and Encrypted Data
During the demand-shaping process, there is a risk of malicious interventions aimed at disrupting the system. In addition, data can be accessed, monitored, or manipulated by malicious parties, raising serious privacy concerns. Such abuses not only threaten system functionality but also highlight the need for robust security measures. The research highlighted in [15] underscores the importance of addressing security vulnerabilities, emphasizing that AI can potentially increase cyber attacks. Another difficulty for GenAI in performing demand-planning operations is its approach to encrypted data. Therefore, while GenAI is promising in processing various types of data, encrypted data traffic requires further research to ensure both efficiency and security in future wireless networks.
IV-D Information Loss and QoS Degradation
@@
A common flaw in data shaping is the oversimplification of information. Although GenAI algorithms generally perform with good accuracy, there is still a risk of losing contextual information during the demand-shaping process. This can result in the omission of important details and nontrivial nuances, leading to incomplete information for the receiver. Consequently, the SLA and QoS of the users will be affected. Therefore, it is crucial to have efficient mechanisms to prevent information loss and maintain the QoS of the users. In scenarios where information loss is unavoidable, data rescheduling is a potential replacement solution for demand-shaping.
IV-E Consistency, Data Privacy, and Transparency
@@
GenAI models can sometimes produce inconsistent outputs for the same content, leading to unreliable or unpredictable results. This is especially critical in situations such as disasters, where models need to operate smoothly and quickly. Therefore, it is essential to develop and continuously improve validation methods to enhance the reliability of GenAI models. Besides, when processing user demands, the security and confidentiality of data are of high importance. Additionally, the black box nature of GenAI, with its lack of transparency about how algorithms work, can make it difficult for users to trust the process.
@@ V Conclusion
Berk Çiloğlu (cilogluberk@gmail.com) is a M.Sc. student of the EMIMEP at Université de Limoges, Limoges, France. His research focuses on applications of 6G and beyond wireless communication technologies, and non-terrestrial networks.
Görkem Berkay Koç (gorkembrkoc@gmail.com) is a M.Sc. student at Ankara Yıldırım Beyazıt University, Ankara, Türkiye. His research interests focus on optimizing wireless network performance, specifically non-terrestrial networks, and visible light communications.
Afsoon Alidadi Shamsabadi (afsoonalidadishamsa@sce.carleton.ca) is a Ph.D. Candidate in the CU-NTN Lab at Carleton University, Canada. Her research focuses on the innovative integration of NTN with terrestrial wireless networks.
Metin Ozturk (metin.ozturk@aybu.edu.tr) is Assistant Professor at Ankara Yıldırım Beyazıt University, Turkiye. His current research includes intelligent networking for wireless communications, with a particular focus on NTN and sustainability.
Halim Yanikomeroglu (halim@sce.carleton.ca) is a Chancellor’s Professor in the Department of Systems and Computer Engineering at Carleton University, Ottawa, Canada, leading the CU-NTN Lab. His research group focuses on wireless access architecture for the 2030s and 2040s, with a particular emphasis on NTN.